OpenStudy (anonymous):

How do you find the sixth term of a geometric sequence with t5 = 24 and t8 = 3.

OpenStudy (anonymous):

t5 = a*r^4 = 24 t8 = a.r^7 = 3 Dividing one by the other r^3 = 1/8 r = 1/2 t6 = a.r^5 = t5*r = 24 * 1/2 = 12

OpenStudy (anonymous):

@heather040200 thank youu :D

OpenStudy (anonymous):

no problem :)